OpenStudy (anonymous):

t-5/12+4 I am normally a good algebra student but for some reason nothing I do with this feels right please help!

OpenStudy (anonymous):

t-5=4*12 t=48+5 t=...?

OpenStudy (angela210793):

then i guess u know how 2 solve it now :):)

OpenStudy (anonymous):

t=53 =D